Circular No. 6589

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ab IN ANONYMOUS GALAXY H.-J. Hagen and D. Reimers, Hamburg Observatory, report the discov-

ery of a supernova (mag 16.2) on a spectrum taken on Feb. 28. SN 1997ab is first visible on an objective-prism plate taken on 1996 Apr. 11 with the Calar Alto 0.80-m Schmidt telescope for the Hamburg Quasar survey; it was automatically selected on the digitized plate as a low-redshift QSO candidate due to its ultraviolet excess and its strong Hβ emission. From the objective-prism spectrum, we derive mag B = 14.7 ± 0.5. A CCD image taken with the Calar Alto 2.2-m telescope shows a bright point source 7′′ southwest of the galaxy center; the supernova is located at α = 9^{h}51^{m}00^{s}.4, δ = +20^{o}04′24′′ (equinox 2000.0). The spectrograms taken with the same instrument show SN 1997ab to be of the rare emission-line Seyfert-1 type II, similar to SN 1987F, with a blue underlying continuum. Besides strong Balmer lines, the spectrum shows broad Fe II emission bumps and a strong infrared Ca II triplet. The redshift of the parent galaxy (from sharp, forbidden H II-region lines O III and S II) is 0.012. SN 1997ab faded in eleven months by < 2 mag and was at least 2 mag brighter in Apr. 1996 than the parent galaxy, which is a dwarf irregular of mag 17.

COMET C/1995 O1 (HALE-BOPP) T. Farnham and D. Schleicher, Lowell Observatory; and S. Lederer,

University of Florida, write: “Narrow-band photometry obtained with the 0.79-m telescope at Lowell Observatory on Mar. 6.5 and 7.5 UT (when r = 1.02–1.01 AU) yield the following mean Haser-model production rates: log Q(OH) = 30.24, log Q(CN) = 28.05, log Q(C_{2}) = 28.16. The dustproduction rate was log (Afρ) = 5.93. Combining these results with those obtained from Lowell by Schleicher, Farnham, and R. Millis during the previous seven weeks at r = 1.54–1.52, 1.34, 1.23–1.20, and 1.09 AU indicates that gas production has now leveled off and dust production is increasing at a slower rate, as compared to the r-dependent slopes of −3 to −4 displayed by both the gas and the dust from mid-Jan. to mid-Feb. However, this leveling off in the gas may in part be due to the rapidly-decreasing geocentric distance yielding smaller projected apertures, coupled with a strong aperture effect exhibited by the various gas species. In contrast, the derived dust production shows little dependence with aperture size.”

Naked-eye m_{1} estimates: Mar. 9.74 UT, −0.3 (K. Cernis, Moletai Observatory, Lithuania); 12.19, −0.4 (M. V. Zanotta, San Marco Pass, Italy); 16.44, −0.5 (G. W. Kronk, Troy, IL); 17.80, −0.6 (P. Candy, Viterbo, Italy).

1997 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green